Spring an a centripetal force free The device is mounted 1.2 ha ticon tally as in lab experiment At acertain speed of rotation, a mass of am tach to the spring extends the Spina by 5 cms. Find speed of rotation revolutions pes min RPM
Answer #1

Given

spring constant k = 1.2 N/m

streche in the spring due to rotation of mass m = 0.075 kg in circular motion , is 0.05 m

the time period is T = 2pi sqrt(m/k) = 2pi sqrt(0.075/1.2) = 1.571 s


   angular frequency w = 2pi/T rad/s


           W = 2pi/(T*2pi) rev/s


           W = 60*2pi/(T*2pi) rpm

           W = 60/ T rpm


           W = 60/1.571 = 38.192 rpm = 38 rpm